OpenStudy (anonymous):

Jacob made a circle-shaped poster for his geometry class. If the radius of circle-shaped poster is 9 inches, what is the approximate circumference?

OpenStudy (anonymous):

I think it would be 28.26

OpenStudy (nincompoop):

let us start by identifying the formula for a circumference (perimeter of a circle) \(C = 2 \pi r \)

OpenStudy (anonymous):

how would i do that

OpenStudy (nincompoop):

since the the unit is provided in inches, we want to be consistent and thus our radius becomes \(r = 9 ~in. \) taking into the consideration of the circumference formula \(C = 2 \pi r \) it follows that our circumference then would be \(C = \pi 2r \rightarrow C= \pi \times 18 ~ in. \) or simply \(C = 18 \pi \)

OpenStudy (anonymous):

So the circumference is 18?

OpenStudy (nincompoop):

no, it IS \(18 \pi \)

OpenStudy (nincompoop):

inches as the unit

OpenStudy (anonymous):

A 56.52 in B 254.34 in C 28.26 in D 113.04 in

OpenStudy (nincompoop):

that is not even an approximation, but rather a very precise measurement :) now if you want to use an approximation, you can use \(\pi = 3.14 \) then it follows that the circumference will be \(C = 18 \pi \rightarrow C= 18~in. \times 3.14 \)

OpenStudy (anonymous):

56.52!

OpenStudy (nincompoop):

but do you understand how it is done now?

OpenStudy (anonymous):

yep! thank you!

OpenStudy (nincompoop):

you begin by identifying the formula for what it is asking for, then plugging in the values